Solve and check the equation: 6y=9x3 Answer: 9=(3/2)x3

Answer by stanbon(75887)

Solve and check the equation: 6y=9x^3
---
y = (9/6)x^3
y = (3/2)x^3
======================
Answer: y = (3/2)x^3
==============
Cheers,
Stan H.
==============